# On the creation of narrow AI: hierarchy and nonlocality of neural network skills

### v1

@@ Abstract
We study the problem of creating strong, yet narrow, AI systems. While recent AI progress has been driven by the training of large general-purpose foundation models, the creation of smaller models specialized for narrow domains could be valuable for both efficiency and safety. In this work, we explore two challenges involved in creating such systems, having to do with basic properties of how neural networks learn and structure their representations. The first challenge regards when it is possible to train narrow models from scratch. Through experiments on a synthetic task, we find that it is sometimes necessary to train networks on a wide distribution of data to learn certain narrow skills within that distribution. This effect arises when skills depend on each other hierarchically, and training on a broad distribution introduces a curriculum which substantially accelerates learning. The second challenge regards how to transfer particular skills from large general models into small specialized models. We find that model skills are often not perfectly localized to a particular set of prunable components. However, we find that methods based on pruning can still outperform distillation. We investigate the use of a regularization objective to align desired skills with prunable components while unlearning unnecessary skills.
@@ 1 Introduction
In this work, we investigate some basic questions about how neural networks learn and represent skills that are relevant to the problem of creating narrow AI systems. As summarized in Figure 1, we focus on two main themes. First, we consider the question of when it is possible to train well-performing neural networks from scratch on a narrow data distribution. Through experiments on a novel synthetic task with hierarchical structure, we find that it can be necessary to train networks on a broad distribution to efficiently learn narrow tasks within that distribution. These results contribute to a growing literature on how task structure influences neural network learning dynamics [12, 13, 14, 15], but with special relevance to the problem of creating narrow AI. Second, we consider the question of whether one can use pruning to turn broad networks into smaller narrow ones. We find that the nonlocality of network representations to prunable model components poses a challenge for this goal. While distributed representations have been extensively studied in the context of neural network interpretability [16] and in the classical connectionist literature [17, 18], we study how this property of neural network computation impacts pruning and unlearning [5] for creating narrow AI. Our specific contributions are as follows:
Figure 1: We study two challenges to making strong, narrow-purpose AI models. (A): Data may have hierarchical structure. If skills have a hierarchical dependence, where some skills are only learnable after more primitive skills are learned first, then it sometimes may be necessary to train on a broad distribution of data to learn certain narrow skills within that distribution. These dynamics may mean that general-purpose models must be trained to achieve certain performance on some domains. (B): Model features are distributed. By default, skills may not be localizable to a particular set of model components (e.g. neurons). In this case, pruning of model components won’t precisely retain wanted skills and remove unwanted skills from models. We explore methods for aligning the model features relevant to particular domains with a smaller subset of model components while unlearning others.
@@
This work is organized as follows. In Section 2, we briefly describe methods. In Section 3 we perform a detailed case study of training and pruning on the CMSP task. We define the task (Section 3.1), study network training dynamics on it (Section 3.2), find distributed representations in these networks (Section 3.3), and use a regularization method for pruning and unlearning in these networks (Section 3.4). We then compare various methods for creating narrow models, studying MNIST in Section 4 and language models in Section 5. We discuss related work in Section 6 and conclude in Section 7.
@@ 3.1 Defining compositional multitask sparse parity (CMSP)
the mean loss exhibits power-law scaling [31, 32] while individual subtasks are learned at different times proportional to their frequency. They conjecture that LLM learning dynamics may be similar.
@@
Figure 2: Training dynamics on compositional multitask sparse parity. Top: training dynamics for a single network trained on four atomic subtasks {0}, {1}, {2}, {3}, and their composition {0,1,2,3}. Bottom left: loss on compositional subtask {0,1,2,3}, when training only on samples from that task vs. also training on atomic subtasks. We see that by training on a broader distribution, we are able to learn the narrow task much faster and more reliably than when only training on samples from the narrow task. Bottom right: When training on the full distribution, deeper networks learn the compositional task faster and more reliably than networks with a single hidden layer. We report the minimum loss within the previous 100 steps of training to filter out loss spikes.
@@ 3.2 Learning dynamics on CMSP
We first train ReLU MLPs with 1-2 hidden layers of width 128 with the Adam optimizer on CMSP samples with $m=4$, $n=64$, $k=4$, and 2000 samples per task (atomic or composite) per batch. We show the learning dynamics of these networks in Figure 2. We find that when we train on a dataset containing both atomic and compositional samples $\mathcal{D}_{\{0\}}\cup\mathcal{D}_{\{1\}}\cup\mathcal{D}_{\{2\}}\cup\mathcal{D}_{\{3\}}\cup\mathcal{D}_{\{0,1,2,3\}}$ in equal proportion, atomic tasks are learned before composite ones.
@@
Something interesting happens however when we remove atomic samples from the dataset: we find that learning composite tasks takes dramatically longer. When we train on $\mathcal{D}_{\{0\}}\cup\mathcal{D}_{\{1\}}\cup\mathcal{D}_{\{2\}}\cup\mathcal{D}_{\{3\}}\cup\mathcal{D}_{\{0,1,2,3\}}$ with 10000 samples per batch (2000 per subtask), across 40 seeds, we find that 27/40 networks converge on composite subtask $\mathcal{D}_{\{0,1,2,3\}}$ within $2\times 10^{9}$ samples, and the networks that do converge typically converge within $2\times 10^{8}$ samples. However, when we train just on $\mathcal{D}_{\{0,1,2,3\}}$, with a batch size of 2000 samples, we find that 0/40 networks converge within $2\times 10^{9}$ samples in Figure 2 (bottom left). It is much more efficient to train on a broader distribution in order to learn subtask $\mathcal{D}_{\{0,1,2,3\}}$ than just training on $\mathcal{D}_{\{0,1,2,3\}}$ on its own.
This result makes sense given the exponential hardness of learning parities [33]. One explanation of these results is that networks compute the parity of the atomic subtask bits in the first hidden layer, and then to learn the composite subtask they compute the parity of these values in the second layer. This way, the network never needs to directly learn the parity of 16 bits, and learning the composite subtask is akin to computing the parity of 4 bits. We find that there is indeed a learning advantage to depth: when we train on $\mathcal{D}_{\{0\}}\cup\mathcal{D}_{\{1\}}\cup\mathcal{D}_{\{2\}}\cup\mathcal{D}_{\{3\}}\cup\mathcal{D}_{\{0,1,2,3\}}$ networks with 2 hidden layers are able to learn the composite task somewhat more reliably (27/40 seeds converged with depth 3 versus 19 with depth 2) but, more significantly, they learn much faster than networks with only a single hidden layer. [^2] Though intriguingly not as slowly as when training on only composite samples, so there is an advantage to training on a broad distribution that does not come from composing the early-layer features in later-layers. In Figure 7, we confirm that this effect is likely not just due to differences in network parameter count. This result may also provide some explanation for advantage of depth in deep learning – not only is depth necessary for networks to efficiently approximate certain functions [34], we find here that depth is helpful to efficiently learn tasks with hierarchical structure.
@@
These sorts of dynamics may be a part of the explanation for why large-scale general-purpose models perform so strongly at many narrow, valuable tasks. We want to emphasize, however, that the toy task where we observe these curriculum effects is fairly contrived. It is not clear to what extent there are similarly strong effects on real-world tasks, and in many domains it is in fact possible to train narrow specialized models by only training on data in that domain. For instance, self-driving cars do not need to be trained on a broad corpus of text like LLMs.
@@ 3.3 Nonlocal representations in CMSP networks
In our CMSP networks, we observe a related problem. By default, without any explicit regularization, there is no incentive for the network to localize certain circuits cleanly into a particular set of neurons. We train networks with two hidden layers on a dataset of CMSP samples with $m=6$, $n=18$, $k=3$ with two different skill trees: {0}, {1}, {2}, {0,1,2}, and {3}, {4}, {5}, {3,4,5} until convergence. In Figure 3, we visualize the connectivity of a 2-hidden-layer MLP trained on this dataset, and see that the network is densely connected without obvious structure.
We attempt to prune this network to retain performance on subtask {0,1,2} while unlearning {3,4,5}. With MLPs, our groups of parameters $g$ are the in-weights, bias, and out-weights for each hidden layer neuron, so that $|G|$ is the total number of hidden neurons. We compute ablation scores on the distribution $\mathcal{D}_{N}=\mathcal{D}_{\{0,1,2\}}$ (estimated on 2000 samples) and prune greedily, as described in Section 2. In Figure 3, we show accuracy on subtask {0,1,2} vs. sparsity with this pruning strategy. When applied naively, network performance tends to degrade quickly, since without explicit regularization the network is not optimized to be naively prunable (see in Appendix Figure 8 and Figure 9 for more on how ablation scores for each subtask vary across neurons). However, when we perform an additional 1000 steps of training on $\mathcal{D}_{\{0,1,2\}}$ (5000 samples per batch) after pruning (after either removing neurons from the architecture or pinning pruned weights at zero) we can recover performance at higher sparsity levels. Since we are training just on compositional samples, 1000 steps is not enough to re-learn this task on its own, so if we can recover performance, that will be because the mechanisms for the task were somewhat preserved after pruning. We observe that often, the tasks {0,1,2} and {3,4,5} are entangled – when we prune as aggressively as we can while being able to recover performance on subtask {0,1,2}, we can often still recover some performance on subtask {3,4,5}. However, for our CMSP networks the degree of entanglement is highly seed-dependent, and sometimes the subtasks are disentangled enough that pruning as aggressively as possible on one subtask does robustly unlearn the other. We show pruning curves across seeds and widths in Appendix Figure 10.
@@
Figure 3: Top: We visualize the connectivity of 2-hidden-layer MLPs trained on a CMSP distribution with subtasks {0}, {1}, {2}, {0,1,2}, {3}, {4}, {5}, {3,4,5}, visualized before (left) and after (right) regularizing network weights with the group lasso sparsity penalty while training on subtask {0,1,2}. We find that network connectivity becomes sparse after regularizing. Bottom: we show how pruning affects task performance on subtasks {0,1,2} and {3,4,5} at varying sparsity levels. We prune neurons based on the absolute change that ablating them has on the loss on subtask {0,1,2}. We find that subtasks here are nonlocal and entangled in the “pretrained” network (left). As we prune neurons according to their relevance on subtask {0,1,2}, at the sparsity at which performance on subtask {0,1,2} accuracy drops below 98% (green line), we can still recover some performance on subtask {3,4,5} with a small amount of additional training. Thus naive pruning here has not completely and robustly unlearned subtask {3,4,5}. However, after regularizing the weights (right), we find that not only we can more aggressively prune the network, but we have also robustly unlearned subtask {3,4,5}. Note that the degree to which subtasks are nonlocal and entangled in the pretrained networks depends on seed and width, and we show a variety of additional curves in Figure 10.
@@ 3.4 Regularizing to “narrow” networks
We find that we can use a simple regularization penalty to simultaneously unlearn some tasks while incentivizing the network to move its features to be less distributed, allowing for more aggressive pruning. As described in Section 2, we simply perform additional training on $\mathcal{D}_{\{0,1,2\}}$ with a group lasso sparsity penalty on the network weights. We aim for this penalty to “clean up” circuitry [41] not relevant to prediction on $\mathcal{D}_{\{0,1,2\}}$ while also sparsifying the weights across groups of parameters, allowing for easier pruning.
We apply this regularization to our CMSP networks, training with penalty strength $\lambda=10^{-3}$ with a batch size of 2000 for 10000 steps. With this regularized network, we then re-compute neuron ablation scores and prune. In Figure 3 we find that, in CMSP networks, this method is effective at unlearning skills and making the network more prunable. We find that we can prune more aggressively while retaining performance on subtask {0,1,2}, and we are not able to recover performance on {3,4,5} at any sparsity level.
@@ 4 Pruning vs distillation: MNIST
Figure 4: Left: We compare the performance of distillation, training from scratch, and two pruning approaches for creating small networks that classify MNIST even digits. Pruning-based approaches Pareto-dominate distillation, achieving high compression ratios with fewer datapoints. All points are averaged over 10 individual training runs. Right: When pruning using group lasso, it often helps to first prune rapidly, degrading performance, and then recover performance with no regularization. Each line represents a single training run, with a new point logged every 2,000 datapoints.
@@
When using distillation (2), we use the approach of Hinton et al. [26] with $T=20$. When pruning with the group lasso penalty (3), we use $\lambda$ values ranging from $0.001$ to $0.008$. Unlike in Section 3.4, we regularize and prune simultaneously, pruning neurons when their L2 norm drops below 0.05. When the number of remaining neurons drops below a target threshold, we remove the pruning penalty and continue training to recover lost performance during pruning. When we prune up front and then separately recover performance (4), we use attribution scores as described in Section 2.
@@ 5 Pruning vs distillation: LLMs on Python documents
We first prune the neurons in the MLP blocks of Llama-3.2-1B [43]. Later, we will also consider pruning residual stream dimensions, which involves pruning all model parameters that “read to” or “write from” a dimension of the residual stream [37]. We use attribution scores when pruning, and show that the attribution scores correlate moderately with true ablation scores in Appendix Figure 11. In Figure 5 (left), we show neuron sparsity vs loss curves. When pruning neurons naively, we see that loss increases quickly at low sparsity levels. We also experiment with applying group lasso regularization while further training on Python documents (learning rate of 2e-6, max length 512, 18 documents per batch) and find that this training does indeed level out the sparsity vs. loss curve, albeit at a slight cost to the loss. Fortunately, we find that we can recover lost performance after pruning by doing a small amount of additional training on Python documents in Figure 5 (right). We find that despite the loss increasing substantially after naively pruning, we can also quickly recover that lost performance, and overall this strategy seems to be more efficient than using group lasso training. We therefore next compare naive pruning + recovery training against distillation and training networks from scratch.
@@
Figure 5: Left: Neuron sparsity vs. loss curve for networks tuned with group lasso regularization with varying $\lambda$ for 70k steps vs. base network. Regularization flattens the sparsity vs. loss curve, at the cost of slightly degrading model performance. Right: after pruning our networks to 30%, 63%, and 80% sparsity for our runs with $\lambda$ of 5e-4, 3e-4, and 1e-3, respectively, we recover performance with additional training. We find that we can recover performance lost during pruning, including in the network that was pruned without first using group-lasso regularization training.
@@
We train networks with the Llama 3 [43] architecture of varying shape and size. We use a learning rate of 5e-4, sequence length of 1024, and batch size of 64. For distillation, we use Llama-3.1-8B as a teacher with $T=2$. For pruning + recovery, we prune Llama-3.2-1B to varying levels of neuron and residual stream sparsity, shown in Appendix Table 2. In Appendix Figure 12 we show learning curves for these three approaches. In Figure 6, we tentatively find that pruning substantially outperforms training from scratch and distilling a model from scratch on the data-parameter frontier. Given a target narrow network size and a fixed data budget, if one already has access to a general model, it appears to be more efficient to prune that model than it is to perform distillation.
We discuss one last finding: pruning random model components performs about as well as pruning lowest-attribution components [44]. In Appendix Figure 13, we find that after a moderate number of recovery steps, attribution pruning and random pruning result in the same recovered performance. This result is in line with our earlier discussion of nonlocality, and the empirical findings of Bricken et al. [16], that monosemantic features are distributed widely across model components. While some studies have found some geometric similarity between functionally similar features [45, 46], in our case it does not seem like the relevant features for our task are localized into a set of “Python” vs. “non-Python” neurons, at least that attribution pruning identifies.

### v2

@@ Abstract
We study the problem of creating strong, yet narrow, AI systems. While recent AI progress has been driven by the training of large general-purpose foundation models, the creation of smaller models specialized for narrow domains could be valuable for both efficiency and safety. In this work, we explore two challenges involved in creating narrow AI systems, having to do with basic properties of how neural networks learn and structure their representations. The first challenge regards when it is possible to train narrow models from scratch. Through experiments on a synthetic task, we find that it is sometimes necessary to train networks on a wide distribution of data to learn certain narrow skills within that distribution. This effect arises when skills depend on each other hierarchically, and training on a broad distribution introduces a curriculum which substantially accelerates learning. The second challenge regards how to transfer particular skills from large general models into small specialized models. We find that model skills are often not perfectly localized to a particular set of prunable components. However, we find that methods based on pruning can still outperform distillation. We investigate the use of a regularization objective to align desired skills with prunable components while unlearning unnecessary skills.
@@ 1 Introduction
In this work, we investigate some basic questions about how neural networks learn and implement skills that are relevant to the problem of creating narrow AI systems. As summarized in Figure˜1, we focus on two main themes. First, we consider the question of when it is possible to train well-performing neural networks from scratch on a narrow data distribution. Through experiments on a novel synthetic task with hierarchical structure, we find that it can be necessary to train networks on a broad distribution to efficiently learn narrow tasks within that distribution. These results contribute to a growing literature on how task structure influences neural network learning dynamics [12, 13, 14, 15], but with special relevance to the problem of creating narrow AI. Second, we consider the question of whether one can use pruning to turn broad networks into smaller narrow ones. We find that the nonlocality of network representations to prunable model components poses a challenge for this goal. While distributed representations have been extensively studied in the context of neural network interpretability [16] and in the classical connectionist literature [17, 18], we study how this property of neural network computation impacts pruning and unlearning [5] for creating narrow AI. Our specific contributions are as follows:
Figure 1: We study two challenges to making strong, narrow-purpose AI models. (A): Data may have hierarchical structure. If skills have a hierarchical dependence, where some skills are only learnable after more primitive skills are learned first, then it sometimes may be necessary to train on a broad distribution of data to learn certain narrow skills within that distribution. These dynamics may mean that general-purpose models must be trained to achieve good performance on some domains. (B): Model features are distributed. By default, skills may not be localizable to a particular set of model components (e.g. neurons). In this case, pruning of model components won’t precisely retain wanted skills and remove unwanted skills from models. We explore methods for aligning the model features relevant to particular domains with a smaller subset of model components while unlearning others.
@@
This work is organized as follows. In Section˜2, we briefly describe methods. In Section˜3 we perform a detailed case study of training and pruning on the CMSP task. We define the task (Section˜3.1), study network training dynamics on it (Section˜3.2), find distributed representations in these networks (Section˜3.3), and use a regularization method for pruning and unlearning in these networks (Section˜3.4). We then compare various methods for creating narrow models, studying MNIST in Section˜4 and language models in Section˜5. We discuss related work in Section˜6 and conclude in Section˜7.
@@ 3.1 Defining compositional multitask sparse parity (CMSP)
the mean loss exhibits power-law scaling [31, 32] while individual subtasks are learned at different times proportional to their frequency [33]. They conjecture that LLM learning dynamics may be similar.
@@
Figure 2: Training dynamics on compositional multitask sparse parity. Top: training dynamics for a single network trained on four atomic subtasks {0}, {1}, {2}, {3}, and their composition {0,1,2,3}. Bottom left: loss on compositional subtask {0,1,2,3}, when training on atomic subtasks and their composition, as in the top subplot, across 40 network seeds (depth 3, width 128). Bottom right: loss on compositional subtask {0,1,2,3}, when only training on samples from that composite task, without also training on atomic tasks, across 40 network seeds (depth 3, width 128). We find that removing the atomic tasks prevents our networks from learning the composite task. We report the minimum loss within the previous 100 steps of training to filter out loss spikes.
@@ 3.2 Learning dynamics on CMSP
We first train ReLU MLPs with 1-2 hidden layers of width 128 with the Adam optimizer on CMSP samples with $m=4$, $n=64$, $k=4$, and 2000 samples per task (atomic or composite) per batch. We show the learning dynamics of these networks in Figure˜2. We find that when we train on a dataset containing both atomic and composite samples $\mathcal{D}_{\{0\}}\cup\mathcal{D}_{\{1\}}\cup\mathcal{D}_{\{2\}}\cup\mathcal{D}_{\{3\}}\cup\mathcal{D}_{\{0,1,2,3\}}$ in equal proportion, atomic tasks are learned before composite ones.
@@
Something interesting happens when we remove atomic samples from the dataset: we find that learning composite tasks takes dramatically longer. When we train on $\mathcal{D}_{\{0\}}\cup\mathcal{D}_{\{1\}}\cup\mathcal{D}_{\{2\}}\cup\mathcal{D}_{\{3\}}\cup\mathcal{D}_{\{0,1,2,3\}}$ with 10000 samples per batch (2000 per subtask), across 40 seeds, we find that 27/40 networks converge on composite subtask $\mathcal{D}_{\{0,1,2,3\}}$ within $2\times 10^{9}$ samples, and the networks that do converge typically converge within $2\times 10^{8}$ samples (bottom left). However, when we train just on $\mathcal{D}_{\{0,1,2,3\}}$, with a batch size of 2000 samples, we find that 0/40 networks converge within $2\times 10^{9}$ samples in Figure˜2 (bottom right). It is much more efficient to train on a broader distribution in order to learn subtask $\mathcal{D}_{\{0,1,2,3\}}$ than just training on $\mathcal{D}_{\{0,1,2,3\}}$ on its own.
This result makes sense given the exponential hardness of learning parities [34]. Since the atomic task bits are disjoint in CMSP, networks can compute the parity of the composite tasks by computing the parity of features learned for the atomic tasks, reducing the complexity of the problem. For instance, one could imagine that depth-3 networks would compute the parity of the atomic tasks in the first layer, then the composite tasks in the next layer. Curiously though, we find that single-hidden-layer (depth-2) networks can still take advantage of a curriculum in Appendix˜A. We leave a mechanistic analysis of how networks learn these tasks to future work.
@@
These sorts of dynamics may be a part of the explanation for why large-scale general-purpose models perform so strongly at many narrow, valuable tasks. We emphasize, however, that the toy task where we observe these curriculum effects is fairly contrived. It is not clear to what extent there are similarly strong effects on real-world tasks, and in many domains it is in fact possible to train narrow specialized models by only training on data in that domain. For instance, self-driving cars do not need to be trained on a broad corpus of text like LLMs.
@@ 3.3 Nonlocal representations in CMSP networks
In our CMSP networks, we observe a related problem. By default, without any explicit regularization, there is no incentive for the network to localize certain circuits cleanly into a particular set of neurons. We train networks with two hidden layers on a dataset of CMSP samples with $m=6$, $n=18$, $k=3$ with two different skill trees: {0}, {1}, {2}, {0,1,2}, and {3}, {4}, {5}, {3,4,5} until convergence. In Figure˜3, we visualize the connectivity of a 2-hidden-layer MLP trained on this dataset, and see that the network is densely connected without obvious structure.
@@
Figure 3: Top: We visualize the connectivity of 2-hidden-layer MLPs trained on a CMSP distribution with subtasks {0}, {1}, {2}, {0,1,2}, {3}, {4}, {5}, {3,4,5}, visualized before (left) and after (right) regularizing network weights with the group lasso sparsity penalty while training on subtask {0,1,2}. We find that network connectivity becomes sparse after regularizing. Negative weights are shown in blue and positive ones in red, with width proportional to the norm of the weight. Bottom: we show how pruning affects task performance on subtasks {0,1,2} and {3,4,5} at varying sparsity levels. We prune neurons based on the absolute change that ablating them has on the loss on subtask {0,1,2}. We find that subtasks here are nonlocal and entangled in the “pretrained” network (left). As we prune neurons according to their relevance on subtask {0,1,2}, at the sparsity at which performance on subtask {0,1,2} accuracy drops below 98% (green line), we can still recover some performance on subtask {3,4,5} with a small amount of additional training. Thus naive pruning here has not completely and robustly unlearned subtask {3,4,5}. However, after regularizing the weights (right), we find that not only we can more aggressively prune the network, but we have also robustly unlearned subtask {3,4,5}. Note that the degree to which subtasks are nonlocal and entangled in the pretrained networks depends on seed and width, and we show a variety of additional curves in Figure˜10.
We attempt to prune this network to retain performance on subtask {0,1,2} while unlearning {3,4,5}. With MLPs, our groups of parameters $g$ are the in-weights, bias, and out-weights for each hidden layer neuron, so that $|G|$ is the total number of hidden neurons. We compute ablation scores on the distribution $\mathcal{D}_{N}=\mathcal{D}_{\{0,1,2\}}$ (estimated on 2000 samples) and prune greedily, as described in Section˜2. In Figure˜3, we show accuracy on subtask {0,1,2} vs. sparsity with this pruning strategy. When applied naively, network performance tends to degrade quickly, since without explicit regularization the network is not optimized to be naively prunable (see in Appendix Figure˜8 and Figure˜9 for more on how ablation scores for each subtask vary across neurons). However, when we perform an additional 1000 steps of training on $\mathcal{D}_{\{0,1,2\}}$ (5000 samples per batch) after pruning (after either removing neurons from the architecture or pinning pruned weights at zero) we can recover performance at higher sparsity levels. Since we are training just on composite samples, 1000 steps is not enough to re-learn this task on its own, so if we can recover performance, that will be because the mechanisms for the task were somewhat preserved after pruning. We observe that often, the tasks {0,1,2} and {3,4,5} are entangled – when we prune as aggressively as we can while being able to recover performance on subtask {0,1,2}, we can often still recover some performance on subtask {3,4,5}. However, for our CMSP networks the degree of entanglement is highly seed-dependent, and sometimes the subtasks are disentangled enough that pruning as aggressively as possible on one subtask does in fact robustly unlearn the other. We show pruning curves across seeds and widths in Appendix Figure˜10.
@@ 3.4 Regularizing to “narrow” networks
We find that we can use a simple regularization penalty to simultaneously unlearn some tasks while incentivizing the network to move its features to be less distributed, allowing for more aggressive pruning. As described in Section˜2, we simply perform additional training on $\mathcal{D}_{\{0,1,2\}}$ with a group lasso sparsity penalty on the network weights. We aim for this penalty to “clean up” circuitry [41] not relevant to prediction on $\mathcal{D}_{\{0,1,2\}}$ while also sparsifying the weights across groups of parameters, allowing for easier pruning.
We apply this regularization to our CMSP networks, training with penalty strength $\lambda=10^{-3}$ with a batch size of 2000 for 10000 steps. With this regularized network, we then re-compute neuron ablation scores and prune. In Figure˜3 we find that, in CMSP networks, this method is effective at unlearning skills and making the network more prunable. We find that we can prune more aggressively while retaining performance on subtask {0,1,2}, and we are not able to recover performance on {3,4,5} at any sparsity level. These results complete our case study on CMSP: in order to efficiently learn compositional tasks, we train on a broad distribution of tasks. Then, we can regularize the network to turn this broad network into a much smaller one which nevertheless retains the broad network’s performance on the target compositional task.
@@ 4 Pruning vs distillation: MNIST
Figure 4: Left: We compare the performance of distillation, training from scratch, and two pruning approaches for creating small networks that classify MNIST even digits. Pruning-based approaches Pareto-dominate distillation, achieving high compression ratios with fewer datapoints. All points are averaged over 10 individual training runs. Right: When pruning using group lasso, it often helps to first prune rapidly, degrading performance, and then recover performance with no regularization. Each line represents a single training run, with a new point logged every 2,000 datapoints. Darker lines correspond to more aggressive pruning runs that attain lower final neuron counts.
@@
When using distillation (2), we use the approach of Hinton et al. [26] with $T=20$. When pruning with the group lasso penalty (3), we use $\lambda$ values ranging from $0.001$ to $0.008$. Unlike in Section˜3.4, we regularize and prune simultaneously, pruning neurons when their L2 norm drops below 0.05. When the number of remaining neurons drops below a target threshold, we remove the pruning penalty and continue training to recover lost performance during pruning. When we prune up front and then separately recover performance (4), we use attribution scores as described in Section˜2.
@@ 5 Pruning vs distillation: LLMs on Python documents
We first prune the neurons in the MLP blocks of Llama-3.2-1B [43]. Later, we will also consider pruning residual stream dimensions, which involves pruning all model parameters that “read to” or “write from” a dimension of the residual stream [37]. We use attribution scores when pruning, and show that the attribution scores correlate moderately with true ablation scores in Appendix Figure˜11. In Figure˜5 (left), we show neuron sparsity vs loss curves. When pruning neurons naively, we see that loss increases quickly at low sparsity levels. We also experiment with applying group lasso regularization while further training on Python documents (learning rate of 2e-6, max length 512, 18 documents per batch) and find that this training does indeed level out the sparsity vs. loss curve, albeit at a slight cost to the loss. Fortunately, we find that we can recover lost performance after pruning by doing a small amount of additional training on Python documents in Figure˜5 (right). We find that despite the loss increasing substantially after naively pruning, we can also quickly recover that lost performance, and overall this strategy seems to be more efficient than using group lasso training. We therefore next compare naive pruning + recovery training against distillation and training networks from scratch.
@@
Figure 5: Left: Neuron sparsity vs. loss curve for LLMs finetuned with group lasso regularization with varying $\lambda$ for 70k steps, as compared to the base network. Regularization flattens the sparsity vs. loss curve, at the cost of slightly degrading model performance. Right: After pruning our networks to 30%, 63%, and 80% sparsity for our runs with $\lambda$ of 5e-4, 3e-4, and 1e-3, respectively, we recover performance with additional training. We find that we can recover performance lost during pruning, including in the network that was pruned without first using group-lasso regularization training. The plot shows an exponential moving average of batch losses, with individual batch losses shaded in back.
@@
We train networks with the Llama 3 [43] architecture of varying shape and size. We use a learning rate of 5e-4, sequence length of 1024, and batch size of 64. For distillation, we use Llama-3.1-8B as a teacher with $T=2$. For pruning + recovery, we prune Llama-3.2-1B to varying levels of neuron and residual stream sparsity, shown in Appendix Table˜3. In Appendix Figure˜12 we show learning curves for these three approaches. In Figure˜6, we tentatively find that pruning substantially outperforms training from scratch and distilling a model from scratch on the data-parameter frontier. Given a target narrow network size and a fixed data budget, if one already has access to a general model, it appears to be more efficient to prune that model than it is to perform distillation.
@@
We also experiment with pruning random model components instead of components selected by attribution score and find that random pruning performs similarly [44]! In Appendix Figure˜13, we find that after a moderate number of recovery steps, attribution pruning and random pruning result in the same recovered performance. This result is in line with our earlier discussion of nonlocality, and the empirical findings of Bricken et al. [16], that monosemantic features are distributed widely across model components. While some studies have found some geometric similarity between functionally similar features [45, 46], in our case it does not seem like the relevant features for our task are localized into a set of “Python” vs. “non-Python” neurons, at least that attribution pruning identifies.
Unlearning: We lastly evaluate whether neuron pruning can robustly unlearn downstream tasks. We consider three pruning methods applied to Llama-3.2-1B at sparsity levels of 30%, 63%, and 80%: (i) random neuron pruning, (ii) attribution-based pruning on Python code, and (iii) models trained with group-lasso regularization on Python documents from Figure˜5. We evaluate each pruned model on four benchmarks: counterfact [47], AI2-ARC (ARC-Easy) [48], WMDP-Bio, and WMDP-Cyber [49] for hazardous knowledge. We found that the base model performed poorly on WMDP-Chem. For each model, we first measure baseline accuracy after pruning, then fine-tune on each benchmark’s training set with batch size 4, sweeping learning rates between 1e-6 and 1e-4 and reporting the best-performing configuration on the test set. In Table˜1, we find that all pruning techniques, including random pruning, robustly unlearn non-Python skills with roughly similar effectiveness, except that at low sparsity, attribution pruning does not achieve much unlearning. This result contrasts somewhat with our findings on CMSP networks in Sections˜3.3 and 3.4, where training with the group lasso penalty led to stronger unlearning than naive pruning.
[…]
Method Spar. counterfact AI2-ARC WMDP Bio WMDP Cyber
[…]